Light of wavelength (3.222x10^-7)m is shined onto a piece of metal with a threshold frequency of (3.96x10^14)hz. How much energy do the ejected electrons have?
Answer #1

Answer – We are given, wavelength , λ = 3.222*10-7 m ,

threshold frequency, νo = 3.96*1014 Hz

We know the energy of ejected electrons is the difference between the energy of the photon and the work function.

We know, work function = h* νo

So, energy of ejected electron = h*c/ λ - h* νo

     = 6.626*10-34J.s * 3.0*108 m.s-1 / 3.222*10-7 m - 6.626*10-34J.s *3.96*1014 Hz

     = 6.17*10-19 J – 2.62*10-19 J

    = 3.54*10-19 J

Energy in eV –

We know,

= 3.54*10-19 J * 6.24*1018

= 2.21 eV.